OpenStudy (liaannne):

In 1 kilogram of pure water, the mass of hydrogen atoms is approximately 112 grams, or 0.112 kilograms. What is the energy equivalent of this mass in joules?

OpenStudy (anonymous):

you are talking about the Einstein formula "E=mc² " ? in this formula doesnt matter if it is hydrogen or any other kind of substance. 0.112 kg = 0.112 kg x( 3x10^8 m/s)^2= 1.0066e+16 joules http://www.1728.org/einstein.htm

OpenStudy (sweetburger):

E=mc^2 where E is energy, m is mass, and c is the speed of light. E=mc^2 shows the relationship between energy and mass.

OpenStudy (liaannne):

Dude you rock!